Suppose the market for rum can be described by the following equations: Demand: P = 10 – Q Supply: P = Q – 4 where P is the price in US dollars per unit and Q is the quantity in thousands of units. Then: 


1
Expert's answer
2021-06-25T10:49:55-0400

"10-Q=Q-4"


"2Q=10+4"


"2Q=14"


"Q=7"


P=$3

Q=7 Thousands of units